gas has temp of 47C and pressure of 2.10⁵ pascal. What pressure would gas have after isochoric boiling it to 77C



1
Expert's answer
2022-04-14T01:44:05-0400

T1 = 47C = 320 K

T2 = 77C = 350 K

P1 / T1 = P2 / T2

P2 = P1 * T2 / T1

P2 = (2 x 105 * 350 K) / 320 K = 2.2 x 105 (Pa)